\begin{thebibliography}{}

\bibitem[\protect\citeauthoryear{Acerbi}{Acerbi}{2002}]{acerbi02}
Acerbi, C. (2002).
\newblock Spectral measures of risk: A coherent representation of subjective
  risk aversion.
\newblock {\em Journal of Banking \& Finance\/}~{\em 26\/}(7), 1505--1518.

\bibitem[\protect\citeauthoryear{Acerbi}{Acerbi}{2003}]{acerbi03}
Acerbi, C. (2003).
\newblock Coherent representations of subjective risk aversion.
\newblock {\em Risk Measures for the 21st Century\/}, 147.

\bibitem[\protect\citeauthoryear{Artzner}{Artzner}{1997}]{artz97}
Artzner, P. (1997).
\newblock Thinking coherently.
\newblock {\em Risk\/}, 68--71.

\bibitem[\protect\citeauthoryear{Artzner, Delbaen, Eber, and Heath}{Artzner
  et~al.}{1999}]{artz99}
Artzner, P., F.~Delbaen, J.-M. Eber, and D.~Heath (1999).
\newblock Coherent measures of risk.
\newblock {\em Mathematical finance\/}~{\em 9\/}(3), 203--228.

\bibitem[\protect\citeauthoryear{Cont}{Cont}{2001}]{cont01}
Cont, R. (2001).
\newblock Empirical properties of asset returns: stylized facts and statistical
  issues.
\newblock {\em Quantitative Finance\/}~{\em 1}, 223--236.

\bibitem[\protect\citeauthoryear{Cotter and Dowd}{Cotter and
  Dowd}{2006}]{cotter06}
Cotter, J. and K.~Dowd (2006).
\newblock Extreme spectral risk measures: an application to futures
  clearinghouse margin requirements.
\newblock {\em Journal of Banking \& Finance\/}~{\em 30\/}(12), 3469--3485.

\bibitem[\protect\citeauthoryear{Delbaen}{Delbaen}{2002}]{del02}
Delbaen, F. (2002).
\newblock Coherent risk measures on general probability spaces.
\newblock In {\em Advances in finance and stochastics}, pp.\  1--37. Springer.

\bibitem[\protect\citeauthoryear{Dowd, Cotter, and Sorwar}{Dowd
  et~al.}{2008}]{dowd08}
Dowd, K., J.~Cotter, and G.~Sorwar (2008).
\newblock Spectral risk measures: properties and limitations.
\newblock {\em Journal of Financial Services Research\/}~{\em 34\/}(1), 61--75.

\bibitem[\protect\citeauthoryear{Gin{\'e} and Nickl}{Gin{\'e} and
  Nickl}{2008}]{gine}
Gin{\'e}, E. and R.~Nickl (2008).
\newblock Uniform central limit theorems for kernel density estimators.
\newblock {\em Probability Theory and Related Fields\/}~{\em 141\/}(3-4),
  333--387.

\bibitem[\protect\citeauthoryear{Henryk and Silvia}{Henryk and
  Silvia}{2008}]{gzyland08}
Henryk, G. and M.~Silvia (2008).
\newblock On a relationship between distorted and spectral risk measures.
\newblock {\em Revista de econom{\'\i}a financiera\/}~{\em 15}, 8--21.

\bibitem[\protect\citeauthoryear{Overbeck}{Overbeck}{2004}]{overbeck04}
Overbeck, L. (2004).
\newblock Spectral capital allocation.
\newblock {\em Capital\/}.

\bibitem[\protect\citeauthoryear{Pyke and Shorack}{Pyke and
  Shorack}{1968}]{pyke68}
Pyke, R. and G.~R. Shorack (1968).
\newblock Weak convergence of a two-sample empirical process and a new approach
  to chernoff-savage theorems.
\newblock {\em The Annals of Mathematical Statistics\/}~{\em 39\/}(3),
  755--771.

\bibitem[\protect\citeauthoryear{Robbins and Siegmund}{Robbins and
  Siegmund}{1972}]{robbins72}
Robbins, H. and D.~Siegmund (1972).
\newblock On the law of the iterated logarithm for maxima and minima.
\newblock In {\em Proc. 6th Berkeley Symp. Math. Stat. Prob}, Volume~3, pp.\
  51--70.

\bibitem[\protect\citeauthoryear{Rosenblatt}{Rosenblatt}{1956}]{rosen56}
Rosenblatt, M. (1956).
\newblock Remarks on some nonparametric estimates of a density function.
\newblock {\em The Annals of Mathematical Statistics\/}, 832--837.

\bibitem[\protect\citeauthoryear{Sen}{Sen}{1978}]{sen78}
Sen, P.~K. (1978).
\newblock An invariance principle for linear combinations of order statistics.
\newblock {\em Zeitschrift f{\"u}r Wahrscheinlichkeitstheorie und Verwandte
  Gebiete\/}~{\em 42\/}(4), 327--340.

\bibitem[\protect\citeauthoryear{Shorack et~al.}{Shorack
  et~al.}{1972}]{shorack72}
Shorack, G.~R. et~al. (1972).
\newblock Functions of order statistics.
\newblock {\em The Annals of Mathematical Statistics\/}~{\em 43\/}(2),
  412--427.

\bibitem[\protect\citeauthoryear{Swanepoel and Van~Graan}{Swanepoel and
  Van~Graan}{2005}]{swane05}
Swanepoel, J.~W. and F.~C. Van~Graan (2005).
\newblock A new kernel distribution function estimator based on a
  non-parametric transformation of the data.
\newblock {\em Scandinavian Journal of Statistics\/}~{\em 32\/}(4), 551--562.

\bibitem[\protect\citeauthoryear{Tsukahara}{Tsukahara}{2013}]{Tsu14}
Tsukahara, H. (2013).
\newblock Estimation of distortion risk measures.
\newblock {\em Journal of Financial Econometrics\/}~{\em 12\/}(1), 213--235.

\bibitem[\protect\citeauthoryear{Wellner}{Wellner}{1977}]{Wellner2}
Wellner, J.~A. (1977).
\newblock A law of the iterated logarithm for functions of order statistics.
\newblock {\em The Annals of Statistics\/}, 481--494.

\bibitem[\protect\citeauthoryear{Wellner et~al.}{Wellner
  et~al.}{1977}]{Wellner1}
Wellner, J.~A. et~al. (1977).
\newblock A glivenko-cantelli theorem and strong laws of large numbers for
  functions of order statistics.
\newblock {\em The Annals of Statistics\/}~{\em 5\/}(3), 473--480.

\bibitem[\protect\citeauthoryear{Winter}{Winter}{1973}]{winter}
Winter, B. (1973).
\newblock Strong uniform consistency of integrals of density estimators.
\newblock {\em Canadian Journal of Statistics\/}~{\em 1\/}(1-2), 247--253.

\bibitem[\protect\citeauthoryear{Zwet}{Zwet}{1980}]{Van80}
Zwet, W. R.~V. (1980, 10).
\newblock A strong law for linear functions of order statistics.
\newblock {\em Ann. Probab.\/}~{\em 8\/}(5), 986--990.

\end{thebibliography}

\documentclass[11pt]{article}
\usepackage{natbib}
\usepackage{url}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage[flushleft]{threeparttable}
\newcommand{\blind}{1}

\newtheorem{defn}{Definition}
\newtheorem{thm}{Theorem}
\newtheorem{rem}{Remark}
\newtheorem{cor}{Corollary}
\begin{document}

\def\spacingset#1{\renewcommand{\baselinestretch}%
{#1}\small\normalsize} \spacingset{1}


\if1\blind
{
  \title{\bf Kernel Based Estimation of Spectral Risk Measures}
  \author{Suparna Biswas and Rituparna Sen\\
    Applied Statistics Division, Indian Statistical Institute, India}
  \maketitle
} \fi

%

\bigskip
\begin{abstract}
Spectral risk measures (SRMs) belongs to the family of coherent risk measures. A natural estimator for the class of spectral risk measures (SRMs) has the form of $L$-statistics. We propose a kernel based estimator of SRM. We investigate the large sample properties of general $L$-statistics based on i.i.d cases and apply them to our kernel based estimator of SRM. We prove that the estimator is strongly consistent and asymptotically normal. We compare the finite sample performance of the kernel based estimator with that of empirical estimator of SRM using Monte Carlo simulation, where appropriate choice of smoothing parameter and the user's coefficient of risk aversion plays an important role. We estimate the exponential SRM of four future indices, namely, Nikkei 225, Dax, FTSE 100 and Hang Seng using our proposed kernel based estimator.
\end{abstract}

{\it Keywords:} Spectral risk measure, Distortion risk, Coherent risk measure, $L$-statistics
\vfill
1. The data that support the findings of this study are openly available at\\ https://www.macrotrends.net and
https://www.nseindia.com

2. The research is supported by grants from the Indian Statistical Institute.
\newpage
\spacingset{1.45}

\section{Introduction}
In financial market, a risk measure is used to determine the amount of capital to be kept in reserve. The purpose of this reserve is to make the risks taken by financial institutions, such as banks and insurance companies, acceptable to the regulator. A risk measure is a mapping that assigns real numbers to the possible outcomes of a random financial quantity, such as an insurance claim or loss of a portfolio. In recent years attention has turned towards convex and coherent risk measures. The concept of coherent risk measure was introduced by \cite{artz97,artz99}. SRMs proposed by \cite{acerbi02,acerbi03}, belongs to the family of coherent risk measure and hence inherit the properties of such measures. SRM is a weighted average of the quantiles of a loss distribution, the weights of which depend on the user's risk aversion. One nice feature of SRMs is that they relate the risk measure to the user's risk aversion (see \cite{dowd08}). In other words, if two users are faced with the same distribution of possible losses, a spectral risk measure indicates that the more risk-averse user faces a higher risk. \cite{acerbi02} suggests that they can be used to set capital requirements or obtain optimal risk-expected return tradeoffs. \cite{overbeck04} discusses how they might be used for capital allocation, and \cite{cotter06} suggest that SRMs could be used by futures clearinghouses to set margin requirements that reflect their corporate risk aversion.

Spectral risk measure is defined below as in \cite{gzyland08}.
\begin{defn}  Let $\phi\in\mathfrak{L}_1([0,1])$ be an admissible risk measure, then the spectral risk measure is defined by \begin{align}\label{S}M_\phi=-\int_0^1\phi(u)Q_udu,\end{align}
where $\phi$ is called the Risk Aversion Function and $Q_u$ is the quantile function.  \end{defn}

The Risk Aversion Function proposed by \cite{cotter06} is
     \begin{align}\label{S1}\phi(u)=\frac{\beta e^{-\beta(1-u)}}{1-e^{-\beta}}\end{align}
where $\beta\in(0, \infty)$ is the user's coefficient of absolute risk aversion. \cite{dowd08} proposed two more Risk Aversion Functions called power spectral risk measures (PSRMs). These are
\begin{equation*} \phi(u)=\left\{ \begin{array}{ll}\gamma(1-u)^{\gamma-1}& \mathrm{for}\quad \gamma<1 \\
\gamma u^{\gamma-1}& \mathrm{for} \quad\gamma>1.\end{array}\right.\end{equation*}

\cite{dowd08} showed certain properties of the Risk Aversion Functions using a small set of alternative loss distributions, namely standard normal, Cauchy, standard uniform, Beta and Gumbel. They observed that SRMs can have some curious and surprising properties, some of which undermine their usefulness for practical risk management. The coefficient of absolute risk aversion $\beta$ plays a important role in spectral risk measures which is similar to the role played by the confidence level in the value at risk and expected shortfall. \cite{cotter06} mentioned that the higher is $\beta$, the more the user cares about higher losses relative to the others. It is seen that if $\phi(u)=\frac{1}{p}1_{0\geq u\geq p}$ then $M_\phi$ is defined as the Expected Shortfall which is a spectral risk measure. It can be inferred that value at risk is not a spectral risk measure as it is not a coherent risk measure.

In the literature we find very little guidance regarding the estimation of SRMs. But we find a lot of literature regarding the estimation of distortion risk measure. \cite{gzyland08} studied the relationship between SRMs and distortion risk measures and proved that SRMs are equivalent to distortion risk pricing measures, or equivalently, spectral risk functions are related to distortion  functions.

\begin{defn}(\cite{Tsu14}) A distortion risk measure is defined as
\begin{align}\label{R}\rho_D=\int_{[0, 1]}Q_udD(u)=\int_\mathcal{R}xdD\circ F(x),\end{align}
where $D$ is a distortion function.
\end{defn}

Comparing (\ref{S}) and (\ref{R}) we get,
\begin{align*}M_\phi=\rho_D \quad \mathrm{iff}\quad D(u)=-\int_0^u\phi(s)ds \quad \forall u.\end{align*} For $\rho_D$ to be coherent, $D$ must be convex.  A distortion risk measure of the form (\ref{R}) suggests a natural estimator which can be written in the form of an $L$-statistic. Suppose we have independent observations $X_1,\ldots, X_n$ and let $X_{n1}\leq\cdots\leq X_{nn}$ be the order statistics. If we replace $F$ by the empirical cdf $\hat{F}_n$ in equation (\ref{R}), then we get a linear function of the order values which we denote as $\hat{\rho}$
   \begin{align*}\hat{\rho}=\sum_{i=1}^nc_{ni}X_{ni},\end{align*} where $c_{ni}=D(i/n)-D((i-1)/n)$.

Various authors have studied and derived the asymptotic properties of $\hat{\rho}$. \cite{shorack72} and \cite{sen78} derived the asymptotic properties of $\hat{\rho}$ for i.i.d. case under different assumptions. \cite{Wellner2} established certain almost sure ``nearly linear" bounds of $\hat{F}_n$ and its left continuous inverse. \cite{Wellner1} established a strengthened version of the Glivenko-Cantelli theorem for the uniform empirical distribution function and used it to establish the asymptotic property of $\hat{\rho}$ for i.i.d. case. \cite{Van80} generalized the results of \cite{Wellner1} and \cite{sen78} considering i.i.d case. According to VanZwet all smoothness conditions on $g$ and $J$ are unnecessary and the pointwise convergence of $J_n$ can be relaxed (for definition of $g$, $J$ and $J_n$ see section 3.2). \cite{Tsu14} established the asymptotic property of $\hat{\rho}$ considering stationary process.

From the previous studies we have observed that standard asymptotic properties are already obtained for $\hat{\rho}$. But there are no such results for estimators of $\rho$ that involve estimators of the distribution function other than the empirical cdf. In this paper our aim is to consider such an estimator and establish its asymptotic properties. Rest of the paper is organized as follows. In section 2 we propose a kernel based estimator of $\rho$. In section 3 and 4 we establish the asymptotic properties and in the Appendix we give the detailed proof of our results. In section 5 we compare the finite sample performance of the kernel based estimator with that of empirical estimator using Monte Carlo simulation. We observe that appropriate choice of the smoothing parameter and the user's coefficient of risk aversion plays an important role in the estimation of kernel based estimator of spectral risk measure. The comparisons are repeated for different values of $n$ (i.e sample size), $\beta$ and four different models. In section 6 we estimate the exponential SRM of four future indices, namely Nikkei 225, Dax, FTSE 100 and Hang Seng based on the daily return data for the period January 2, 2009 to January 2, 2019 using our proposed kernel based estimator. For comparison purposes, we also present the results for an earlier period namely January 1, 1991 to December 31, 2003 since this is the period considered by \cite{cotter06}. Finally in section 7 we discuss the findings.

\section{Proposed Estimator}
The kernel method introduced by \cite{rosen56} has received considerable attention in nonparametric estimation.  If $X_1,\ldots,\ X_n$ are i.i.d. random variables. Then the usual kernel distribution function is defined as follows
\begin{eqnarray*}F_{n,b}(x)&=&\frac{1}{nb}\sum_{i=1}^n\int_{-\infty}^xk\left(\frac{t-X_i}{b}\right)dt\\
                           &=&\frac{1}{n}\sum_{i=1}^nK\left(\frac{x-X_i}{b}\right).
\end{eqnarray*}
Let us consider the following assumptions:

\emph{Assumption 1:} The kernel distribution function $K$ is differentiable, with a bounded kernel density $k$ with zero mean and finite variance.

\emph{Assumption 2:} $b$ is the smoothing parameter satisfying the condition $b\rightarrow0$ and $nb\rightarrow\infty$ as $n\rightarrow\infty$.

The optimal smoothing parameter suggested by \cite{swane05} is given as follows
$$b=\left[\frac{375\sqrt{3}}{28\pi}\right]^{1/7}\sigma^{-4/7}n^{-1/7},$$
where $\sigma=\min\{S,IQR/1.349\},\ S$ and $IQR$ are the sample standard deviation and inter quartile range respectively.

Based on the usual kernel distribution function i.e. $F_{n,b}(x)$, we propose the following estimator for $\rho$.
\begin{align}\label{B}\hat\rho^b_D=\int_{[0, 1]}F_{n,b}^{-1}(u)dD(u)=\int_\mathcal{R}xdD\circ F_{n,b}(x).\end{align}

\section{Consistency}

We establish the consistency of $\hat\rho^b_D$ by following the techniques used by \cite{shorack72} and \cite{Wellner1}. Let $\xi_1, \xi_2, \ldots$ be a sequence of independent and identically distributed uniform ($0,\ 1$) random variables with distribution function $F (F(t)=t)$ on [$0,\ 1$] and let $F_{n,b}$ denote the kernel distribution function estimator defined as follows. $$F_{n,b}(t)=\frac{1}{nb}\sum_{i=1}^n\int_{-\infty}^tk\left(\frac{x-\xi_i}{b}\right)dx,\ 0\leq t\leq1$$ where $k$ is the kernel density estimator and $b$ is the smoothing parameter satisfying the condition $b\rightarrow0$ and $nb\rightarrow\infty$ as $n\rightarrow\infty$.

\subsection{Convergence of $F_{n,b}$ to $F$}

The convergence of $\hat{F}_{n}-F$, with respect to $d_h$-metric is an important tool in the study of linear rank statistics (\cite{pyke68}) and linear combinations of order statistics (\cite{shorack72}). Similarly, the convergence of $F_{n,b}-F$ with respect to $d_h$-metric is an important tool in our analysis.

 If $h$ is a nonnegative function approaching zero at the endpoints of the interval [$0,\ 1$], and $x, y$ are functions on [$0,\ 1$], the $d_h$-metric is defined by
$d_h(x,y)=d(x/h,y/h)=\sup_{0<t<1}|x(t)-y(t)|/h(t)$, where $d$ denotes the usual supremum metric. Now using the above definition we can define
$$d_h(F_{n,b},F)=d(F_{n,b}/h,F/h)=\sup_{0\leq t\leq1}|F_{n,b}(t)-t|/h(t).$$
From \cite{winter}, we have \begin{align*}d(F_{n,b},F)=\sup_{0\leq t\leq1}|F_{n,b}(t)-t|\rightarrow0 \ \mathrm{as} \ n\rightarrow\infty. \end{align*}

 In Theorem 1 below we establish that $\int_0^1(1/h)dF<\infty$ is both necessary and sufficient for $d_h(F_{n,b},F)\rightarrow0$ with probability one as $n\rightarrow\infty$. Here $\int\cdot\ dF$ denotes integration with respect to Lebesgue measure. Our main motive to establish this type of result is to provide strong laws of large numbers for linear functions of order statistics.

\begin{defn} Let $\mathcal{H}(\nearrow)$ denote the set of all nonnegative, nondecreasing, continuous functions $h$ on [$0,\ 1$] for which $\int_0^1(1/h)dF<\infty$. Let $\mathcal{H}$ denote the set of all functions $h$ such that $h(t)=h(1-t)=\overline{h}(t)$ for $0\leq t \leq1/2$ and some $\overline{h}$ in $\mathcal{H}(\nearrow)$. \end{defn}

\begin{thm} Under Assumptions 1 and 2,

(i) If $h\in\mathcal{H}(\nearrow)$ then \begin{align}\label{L}\lim_{n\rightarrow\infty}d_h(F_{n,b},F)=0 \  w.p.1.\end{align}
(ii) If $h$ is increasing on [0, 1] and $\int_0^1(1/h)dF=+\infty$ then \begin{align}\limsup_{n\rightarrow\infty}d_h(F_{n,b},0)=+\infty\ w.p.1.\end{align} \end{thm}

\begin{rem} (i) of the above Theorem may be extended, using symmetry, to the following $$\lim_{n\rightarrow\infty}d_h(F_{n,b},F)=\lim_{n\rightarrow\infty}d_h(F_{n,b}-F,0)=0 \quad w.p.1 for h\in\mathcal{H}$$. Also, (i) implies that $\lim_{n\rightarrow\infty}d_h(F_{n,b},0)=d_h(F,0) w.p.1 for h\in\mathcal{H}(\nearrow)$.
\end{rem}

\begin{rem} (ii) For $h\in\mathcal{H}(\nearrow)$, we define a process $X_i$ on [$0,1$] where $X_i(t)=\frac{K((t-\xi_i)/b)}{h(t)}$ and write $||l||=d(l,0)$ for $l\in G'[0, 1]\equiv G'$ where $G'[0, 1]$ is the set of right continuous functions on [$0, 1$] with left limits. Then ($G', ||\cdot||$) is an (inseparable) Banach space, and (i) of Theorem 1 is a strong law of large numbers for Banach space valued random elements: $E(X_1)=\frac{F+b^2f^\prime\mu_2(k)/2+o(b^2)}{h}=F/h$, $||X_1||=d_h(K((t-\xi_1)/b),0)=K(0)/h(\xi_1)$, and (i) of Theorem 1 asserts that if $E||X_1||=K(0)\int_0^1(1/h)dF<\infty,$ then
$$\lim_{n\rightarrow\infty}||1/n\sum_1^nX_i-E(X_1)||=0\ w.p.1.$$
\end{rem}
We state a Corollary which is similar to Corollary 1 of \cite{Wellner1}.

\begin{cor}
If $h\in\mathcal{H}(\nearrow)$ then for all $\tau>1$ $$P(d_h(F_{n,b},0)>\tau d_h(F,0)\ for\ some\ 0<t\leq1\ i.o.)=0.$$
\end{cor}

 \cite{Wellner2} proved certain almost sure ``nearly linear" bounds for the empirical distribution function $\hat{F}_n(t)=\frac{1}{n}\sum_{i=1}^n1_{[0, t]}(\xi_i)$ for $0\leq t\leq1$ and $\hat{F}_n^{-1}$, the left continuous inverse of $\hat{F}_n$. In Theorem 2 below, we derive these bounds for $F_{n,b}.$

\begin{thm} Let Assumptions 1 and 2 hold. Let $\tau_1,\tau_2>1$ be fixed. Then there exists $0<\lambda=\lambda(\tau_1,\tau_2)<1/2$ and a set $A\subset\Omega$ with $P(A)=1$ having the following properties: for all $\omega\in A$ there is an $N\equiv N(\omega,\tau_1,\tau_2)$ for which $n\geq N$ implies\end{thm}
\begin{enumerate}
\item $1-\left(\frac{1-t}{\lambda}\right)^{1/\tau_2}\leq F_{n,b}(t)\leq (t/\lambda)^{1/\tau_1}$ for $0\leq t\leq1$,
\item $\lambda t^{\tau_1}\leq  F_{n,b}$ for all $t$ such that $0< F_{n,b}$,
\item $ F_{n,b}\leq1-\lambda(1-t)^{\tau_2}$ for all $t$ such that $ F_{n,b}<1$,
\item $\lambda t^{\tau_1}\leq  F_{n,b}^{-1}(t)\leq1-\lambda(1-t)^{\tau_2}$ for $0\leq t\leq1$,
\item $ F_{n,b}^{-1}(t)\leq (t/\lambda)^{1/\tau_1}$ for $t\geq\frac{1}{n}$, and
\item $1-\left(\frac{1-t}{\lambda}\right)^{1/\tau_2}\leq  F_{n,b}^{-1}(t)$ for $t\leq1-\frac{1}{n}$.
\end{enumerate}

Theorem 1 and Theorem 2 play an important role in establishing a strong law for $T_n$ in next section.

\subsection{Consistency of $\hat{\rho}_D^b$}

We have already observed that a natural estimator for distortion risk measure has the form of $L$-statistics. Let $\mathcal{G}$ denote the set of left continuous functions on ($0,\ 1$) that are of bounded variation on ($\theta,\ 1-\theta$), for all $\theta\in(0,1/2)$; fix $g\in\mathcal{G}$. Let $c_{n1},\ldots,\ c_{nn}$ for $n\geq1$, be known constants. Now, for $0\leq t\leq1$, we define $\psi_n(t)=-\int_t^1J_ndF$ so that $\frac{c_{ni}}{n}=\left[\psi_n\left(\frac{i}{n}\right)-\psi_n\left(\frac{(i-1)}{n}\right)\right]$ for $0\leq t\leq1$.

\begin{eqnarray*}T_n&=&\int_0^1g(F_{n,b}^{-1})J_ndF \\
&=&\frac{1}{n}\sum_{i=1}^ng(\xi_{ni})c_{ni}
\end{eqnarray*}
where $0\leq\xi_{n1}\leq\ldots\leq\xi_{nn}\leq1$ denote the order statistics of the first $n$ i.i.d uniform ($0,\ 1$) random variables.

\begin{rem} If $g=f(I^{-1}),\ f\in\mathcal{G}$ for some distribution function $I$, then $T_n$ has the same distribution as $S_n=\frac{1}{n}\sum_{i=1}^nc_{ni}f(X_{ni}),$ where $X_{n1}\leq\ldots\leq X_{nn}$ are the order statistics of a sample of size $n$ from $I$.\end{rem}

For $n\geq1$, let us define functions $J_n$ on [$0,\ 1$] by $J_n(t)=c_{ni}$ for $(i-1)/n<t\leq i/n$, where $1\leq i\leq n$ and $J_n(0)=c_{n1}$ and set $$\mu_n=\int_0^1gJ_ndF.$$
Now in order to prove our two important result, we define a certain function and assume certain properties drawing upon \cite{shorack72}. For fixed $b_1\ ,b_2>0$ and $M>0$ define a ``scores bounding function" $B$ by
$$B(t)=Mt^{-b_1}(1-t)^{-b_2},\ 0<t<1.$$
For $\delta>0$ define
\begin{eqnarray*}D(t)&=&Mt^{-1+b_1+\delta}(1-t)^{-1+b_2+\delta},\ 0<t<1,\\
                 h(t)&=&[t(1-t)]^{1-\delta/2},\ 0<t<1,\\
                 h^*(t)&=&[t(1-t)]^{1-\delta/4},\ 0<t<1. \end{eqnarray*}
Now, let $g$ be a fixed function in $\mathcal{G}$. Let us denote $J$ to be a fixed measurable function on ($0,\ 1$) and set
\begin{align}\label{MU}\mu=\int_0^1JgdF. \end{align}\\
\emph{Assumption(3):(Boundedness)}. Let $|g|\leq D$, $|J_n|\leq B$ and $|J|\leq B$ on ($0,\ 1$) and suppose that $\int_0^1Bhd|g|<\infty$.\

 \emph{Assumption(4):(Smoothness)}. Except on a set of $t$'s of $|g|$-measure $0$ we have both $J$ is continuous at $t$ and $J_n\rightarrow J$ uniformly in some small neighbourhood of $t$ as $n\rightarrow\infty$.
\begin{thm} If Assumptions 1-3 hold, then\end{thm}
$$\lim_{n\rightarrow\infty}(T_n-\mu_n)=0\ w.p.1.$$

 If $J$ and $g$ satisfy Assumption 3  then $|\mu|<\infty$. We state a Corollary which is similar to Corollary 2 of \cite{Wellner1}.

\begin{cor}If $\lim_{n\rightarrow\infty}\mu_n=\mu_{\infty}$ exists (with $|\mu_{\infty}|<\infty$) and Assumption 3 holds, then $$\lim_{n\rightarrow\infty}T_n=\mu_{\infty}\ w.p.1.$$ \end{cor}

\begin{thm} If Assumptions 1-4 hold, then $$\lim_{n\rightarrow\infty}T_n=\mu\ w.p.1$$
where $\mu$ is finite.\end{thm}

\begin{rem}
From Theorem 4 we can say that our estimator $\hat{\rho}_D^b$ in (\ref{B}) proves to possess strong consistency under the very general conditions stated above.
\end{rem}

\section{Asymptotic Normality}

In this section we establish the asymptotic normality of $\hat{\rho}_D^b$. The technique is similar to \cite{shorack72} and \cite{Tsu14}. \cite{gine} has established the uniform central limit theorems for kernel density estimators. Using Corollary 2 of Section 4 in \cite{gine} we have proved the asymptotic normality of $\hat{\rho}_D^b$. The Corollary is stated below

\begin{cor} (\cite{gine}) Let the random variables $X_1,\ldots, X_n$ are i.i.d. according to the law $\mathbb{P}$ on $\mathbb{R}$ and $\mathbb{B}$ be the Borel-$\sigma$-algebra. Let $C(\mathbb{R})$ denote the Banach space of bounded real-valued continuous functions on $\mathbb{R}$ normed by the usual sup-norm $||\cdot||_\infty$. The variables $X_i$ are taken to be the coordinate projections of the infinite product probability space ($\mathbb{R}^{\mathbb{N}}, \mathbb{B}_{\mathbb{R}^{\mathbb{N}}}, \mathbb{P}^{\mathbb{N}}$). Let $k$ be a kernel of order $r>q+1/2$. Choose $b>0$ of order $b\simeq n^{-1/(2q+1)}$, $q>0$. Then $$\sqrt{n}(F_{n,b}-F)\rightarrow \mathbb{V},$$ where $\mathbb{V}$ is the $\mathbb{P}$-Brownian bridge in $C(\mathbb{R})$. That is \{$\mathbb{V}(t):0\leq t\leq1$\} is a Gaussian process with zero mean and covariance function $\sigma(s,t)=E\mathbb{V}(s)\mathbb{V}(t)=s\wedge t-st$.
\end{cor}

\begin{thm}Let Assumptions 3 and 4 hold. Let $k$ be a kernel of order $r>q+1/2$, for real $q>0$. Choose $b>0$ of order $b\simeq n^{-1/(2q+1)}$. Then $$\sqrt{n}(T_n-\mu_n)\xrightarrow{\text{d}} N(0,\sigma^2),$$ where $$\sigma^2=\int_0^1\int_0^1(s\wedge t-st)J(s)J(t)dg(s)dg(t)<\infty.$$\end{thm}

 From Theorem 5 we have the following result.

\begin{cor}If Assumptions 3 and 4 hold, then $$\sqrt{n}(\hat{\rho}_D^b-\rho)\xrightarrow{\text{d}}N(0,\sigma^2),$$ where $$\sigma^2=\int_0^1\int_0^1(s\wedge t-st)J(s)J(t)dg(s)dg(t)<\infty.$$
\end{cor}

\section{Simulation}
We compare the mean squared error (MSE) of the two estimators of distortion risk measure, viz. the empirical estimator $\hat{\rho}$ and the kernel based estimator using usual kernel distribution function $\hat\rho^b_D$. It is difficult to compute the exact value of the MSE of these estimators even if the the data generating process is completely specified. Therefore we use Monte-Carlo (MC) simulation to approximate the MSE of each of these estimators. The Monte-Carlo (MC) estimate of the MSE of any estimator $P_n$ of a parameter $\theta$ is defined as $\frac{1}{B}\sum^B_{j=1}(P_{nj}-\theta)^2$, where $B$ is the number of MC samples each of size $n$ drawn from a given process and $P_{nj}$ is the estimate based on the $j$th MC sample, $j=1,\cdots,B$. We consider three models.
\begin{eqnarray*}\ & &(i)\ \{X_i\}_{i=1,2,\cdots} \text{is an i.i.d. process, marginal distribution GPD with}\ \xi=1/3.\\
                                           \ & &(ii)\ \{X_i\}_{i=1,2,\cdots} \text{is an i.i.d. process, marginal distribution student's-t with 4 df}.\\
                                           \ & &(iii)\ \{X_i\}_{i=1,2,\cdots} \text{is an i.i.d. process, marginal distribution} N(0,1).                                         \end{eqnarray*}
The first two models are motivated by empirical observations by \cite{cont01} regarding the extent of tail heaviness of the marginal asset return distributions. \cite{cont01} mentioned that when sample moments based on asset return data are plotted against sample size, the sample variance seems to stabilize with increase in sample size. But the behavior of the fourth order sample moment seems to be erratic as $n$ is increased. This feature is also exhibited by the sample moments based on i.i.d. draws from the Student's t distribution with four degrees of freedom, which displays a tail behavior similar to many asset return distributions. Cont also mentioned that the daily return distributions  of stocks, market indices and exchange rates  seem to exhibit power law tail with exponent $\alpha$ satisfying, $\xi=1/\alpha$  varying  between 0.2 and 0.4.\

To study the effect of dependence on the above mentioned estimators of distortion risk measure we consider the following GARCH(1,1) model
\begin{eqnarray*}&&\ (iv)\ X_i=\sigma_iZ_i,\\
                 &&  \sigma^2_i=0.061 X^2_{i-1}+ 0.932\sigma^2_{i-1}.
  \end{eqnarray*}
The model (iv) is the GARCH model fitted to the Nifty 50 daily loss data for the duration 1st January 2009 to 1st January 2019. The data is collected from national stock exchange (NSE) website (https://www.nseindia.com). There are $2476$ daily log return values (log returns are calculated considering the closing value of the index) in our data.

From each of the above models $(i)-(iv)$ and for each combination of $n$ and $\beta$, we draw 1000 MC samples of size $n$. From each of these samples we compute the values of the two estimators of $\rho_D$. From these values we compute the MC estimate of the MSE of that estimator for different choices of $n$, $\beta$ and the underlying model. In each case, let the MC estimates of the MSE of the estimators $\hat{\rho}$ and $\hat\rho^b_D$ be denoted by MSE1 and MSE2 respectively. In Table 1 we report the ratio $\frac{MSE2}{MSE1}$ for $\beta$=1, 5 and 10 and for $n$=30, 100 and 250 considering the four models. The bandwidth chosen is defined in section $2$. We next summarize our findings.

 We observe that in all the cases presented in Table 1, kernel based estimator do better than the empirical estimator $\hat \rho$ for appropriate choice of the smoothing parameter and the user's coefficient of risk aversion $\beta$. For example, for $n=30$ and $\beta=5$ the potential reduction in the MSE ranges between $6$ and $21\%$ for the different models considered above. For $n=100$ and $\beta=5$ the potential reduction in the MSE ranges between $2$ and $6\%$ for the different models. The difference is higher for low sample size and lower values of beta.

\section{Data Analysis}

Our data set consists of daily log returns calculated considering the daily closing prices, for four heavily traded index futures between January 1, 1991 and Jan 2, 2019. The four instruments considered are  FTSE100, DAX, Hang Seng and Nikkei225 futures. We divided the total time into three periods. The first period is between January 1, 1991 and December 31, 2003. With $3280$ daily log return values, this is similar to the data set considered by \cite{cotter06}.   The second period is between January 2, 2004 to December 31, 2008.  With $1250$ daily log return values, this period includes the global financial crisis. The third period is the most recent 10 years between January 2, 2009 to January 2, 2019 with $2582$ daily log return values. These data are collected from Macrotrends website (https://www.macrotrends.net).

The Financial Times Stock Exchange 100 Index, also called the FTSE 100 Index is a  share index of the 100 companies listed on the London Stock Exchange with the highest market capitalisation. It is seen as a gauge of prosperity for businesses regulated by UK company law. The DAX is a blue chip stock market index consisting of the 30 major German companies trading on the Frankfurt Stock Exchange. Prices are taken from the Xetra trading venue. The Hang Seng Index is a freefloat-adjusted market-capitalization-weighted stock-market index in Hong Kong. It is used to record and monitor daily changes of the largest companies of the Hong Kong stock market and is the main indicator of the overall market performance in Hong Kong. These 50 constituent companies represent about $58\%$ of the capitalisation of the Hong Kong Stock Exchange. The Nikkei 225, more commonly called the Nikkei, the Nikkei index, or the Nikkei Stock Average, is a stock market index for the Tokyo Stock Exchange (TSE). It has been calculated daily by the Nihon Keizai Shinbun (The Nikkei) newspaper since 1950. It is a price-weighted index, operating in the Japanese Yen (JP$\yen$), and its components are reviewed once a year.\

We apply the kernel based estimator and estimate the exponential SRM of the FTSE100, DAX, Hang Seng and Nikkei225 futures index, for three periods. We have also estimated the standard deviation and confidence interval. We run $10,000$ sets of Monte Carlo simulation. We have obtained 10,000 estimates of SRM from the simulations and then we have calculated the standard deviation and confidence interval. In Table 2 we have reported the exponential SRM and the standard deviation of the data set for the three periods separately using the $\beta$ values $1$, $5$, $10$, $20$, $100$ and $200$. The main motive of using the different $\beta$ values is to compare the findings with that of the findings of \cite{cotter06}.\ In Table 3 we have reported the corresponding $90\%$ confidence intervals of the SRM estimates.

From Table 2 we observe that FTSE100 is the least risky index and Hang Seng is the most risky index for the first period. This agrees with the results of \ref{cotter06} for the same period. For the other two periods, we observe that while FTSE100 is still the least risky index, Nikkei225 is the most risky index. It is observed that risk is high and so in the variance of the risk, in the second period compared to first and third periods across all the securities and all values of risk aversion. This can be attributed to the global financial crisis that is included in the second period. In fact, the risk is lowest in the third period. We also observe that if we estimate the $90\%$ confidence intervals as described in \citep{cotter06} we obtain similar type of results.

\section{Conclusion}

In this paper we have discussed about the SRMs and their equivalence relation with distortion risk measure. We have discussed and proposed a kernel based estimator of SRM. We have derived certain asymptotic properties of the kernel based estimator of SRM, which has the form of $L$-statistics. The asymptotic results are based on i.i.d. case. The kernel based estimator is strongly consistent and asymptotically normally distributed. We have also derived certain almost sure nearly linear bounds of the kernel distribution function which plays an important role in establishing the strong consistency of the kernel based estimator of SRM.\

From the simulation study it is observed that the choice of the bandwidth and the choice of the absolute risk aversion coefficient plays an important role. We observe that for small sample size ($n\leq250$) and with a preferable choice of absolute risk aversion coefficient $\beta$, the kernel based estimator $\hat{\rho}^b_D$ outperforms the empirical estimator $\hat{\rho}$. It is also seen that the kernel based estimator outperforms the empirical estimator in both i.i.d. and dependent cases. Finally based on our simulation study where we have observed that $\hat{\rho}^b_D$ outperforms $\hat{\rho}$, we estimate the exponential SRM of four heavily traded index futures that is, the FTSE100, DAX, Hang Seng and Nikkei225 futures considering the period from January 1, 1991 to January 2, 2019. The SRM estimates suggest that the FTSE100 is the least risky index and Hang Seng is the most risky index during the period from January 1, 1991 to December 31, 2003. Similar observation can also be seen in \citep{cotter06}, where the authors have estimated the extreme spectral risk measures using the Peaks-over-threshold approach. For the other periods, we find that while FTSE100 is still the least risky index and Nikkei225 is the most risky index. It is also seen that the risk is maximum in the second period across all the securities and all the values of user's coefficient of risk aversion and so is the variance.\\
\newpage
\section*{Appendix}
\begin{defn} Let $X$ be a random variable representing a loss of some financial position and $F$ be the distribution function of $X$. Then, the quantile function is $Q_u=inf\{x:F(x)\geq1-u\},\ 0<u<1$.\end{defn}

\begin{defn}(Delbaen \citep{del02}) A risk measure $\rho$ is a mapping from $\psi$ to $\mathbb{R}$ satisfying certain properties, viz.
\begin{enumerate}
\item $X\geq0\Rightarrow\rho(X)\leq0$.
\item $X\geq Y\Rightarrow\rho(X)\leq\rho(Y),\  X,Y\in \psi.$
\item $\rho(\lambda X)=\lambda\rho(X)$, $\forall\lambda\geq0,\ X\in \psi.$
\item $\rho(X+k)=\rho(X)-k$, $\forall k\in\mathbb{R},\ X\in \psi.$
\end{enumerate}
\end{defn}
The term ``coherent" risk measure is reserved for risk measures that satisfies one more additional property, viz. subadditivity. Artzner et al. introduced the concept of coherent risk measure (see \citep{artz97}, \citep{artz99}).
\begin{defn}(Delbaen \citep{del02})  A risk measure $\rho$ on $\psi$  is said to be coherent if in addition to the properties $1-4$, $\rho$ also satisfies the following ``subadditivity" property, viz.
$$ \rho(X+Y)\leq\rho(X)+\rho(Y), \forall X,\ Y \in \psi.$$
\end{defn}

\begin{defn} (Gzyl and Mayoral \citep{gzyland08}) An element $\phi\in\mathfrak{L}_1([0,1])$ is called an admissible risk spectrum if
\begin{enumerate}
\item $\phi\geq0$
\item $\int_0^1|\phi(t)|dt=1$
\item $\phi$ is non-increasing.
\end{enumerate}
\end{defn}

\begin{defn} (Gzyl and Mayoral \citep{gzyland08}) A function $D:[0,1]\rightarrow[0,1]$ is a distortion function if
\begin{enumerate}
\item D(0)=0 and D(1)=1.
\item D is non-decreasing function.
\end{enumerate}
\end{defn}

\textbf{Proof of Theorem 1.} First we begin with (ii).
Suppose that $h$ is increasing on [$0,\ 1$] and $\int_0^1(1/h)dF=+\infty$.
Now,\begin{eqnarray*} d_F(F_{n,b},0)&=&\sup_{0\leq t\leq1}(F_{n,b}(t)/t)\\
&\geq&F_{n,b}(\xi_{n1})/\xi_{n1}\\
&=&\frac{1}{n\xi_{n1}}\sum_{i=1}^nK\left(\frac{\xi_{n1}-\xi_i}{b}\right) \end{eqnarray*}
Now, from (i) of Theorem 1 of Robbins and Siegmund \citep{robbins72} we see that if $\xi_1, \xi_2,\ldots$ are independent and uniform on ($0,\ 1$) and $V_n=min(\xi_1,\ldots,\xi_n)=\xi_{n1}$. If $c_n=1/r$ for fixed $r$, where $r$ is any arbitary. Now, as $c_n/n\downarrow$ for all sufficiently large $n$ and $\sum_{n=1}^\infty\frac{c_n}{n}$ diverges then $P(nV_n\leq c_n \ i.o.)=1$.
Hence, we can write
\begin{align*}\limsup_{n\rightarrow\infty}d_F(F_{n,b}, 0)=+\infty \ w.p.1. \end{align*}
Now, if $h\leq aF$, for some $a>0$ and using equation (\ref{L}) we have
\begin{eqnarray*}\limsup_{n\rightarrow\infty}d_h(F_{n,b},0)&=&\limsup_{n\rightarrow\infty}\left(\sup_{0\leq t\leq1}\left(\frac{F_{n,b}(t)}{h(t)}\right)\right) \\
&\geq&\limsup_{n\rightarrow\infty}\left(\sup_{0\leq t\leq1}\left(\frac{F_{n,b}(t)}{at}\right)\right) \\
&=&\frac{1}{a}\limsup_{n\rightarrow\infty}d_F(F_{n,b},0) \\
&=&+\infty \ w.p.1. \end{eqnarray*}

If $h\leq aF$ for some $a>0$ does not hold, then for every $a>0$, $h(t)>at$, for some $t\in [0,\ 1]$. Hence, by monotonicity of $h$ this implies that $h\geq aF$, for some $a>0$. \

 Now, let $R_i(t)=\frac{1}{b}\int_{-\infty}^tk\left(\frac{x-\xi_i}{b}\right)dx$ so that $$F_n=\frac{1}{n}\sum_{i=1}^nR_i(t).$$
Let $M>0$ and define events $B_n$ and $D_n$ by
$$B_n=\{d_h(F_{n,b},0)>M\}=\left\{d_h\left(\sum_{i=1}^nR_i,0\right)>nM\right\}$$ and $$D_n=\{d_h(R_n,0)>nM\}$$
Now, since $\sum_{i=1}^nR_i\geq R_n$, $d_{h}\left(\sum_{i=1}^nR_i,0\right)\geq d_h(R_n,0)$ and hence we can write $\{D_n \ i\ . \ o\ .\}\subset\{B_n \ i\ .\ o\ .\}$.
But the events $D_n$ are independent and therefore by Borel-Cantelli Lemma, we have

\begin{align}\label{P}P(D_n\ i\ .\ o\ .)=0\ or\ 1\ according\ as \sum_{n=1}^\infty P(D_n)<\infty\ or\ =\infty\end{align}
Now, we compute $P(D_n)$. Since the $R_i$'s are independent and identically distributed we may drop the subscript $n$; hence for $n$ sufficiently large.
\begin{eqnarray*}P(D_n)&=&P(d_h(R, 0)>nM)\\
&=&P\left(\frac{K(0)}{h(\xi)}>nM\right)\ where,\ K(0)=\int_{-\infty}^0k(u)du\\
&=&P(h(\xi)<K(0)n^{-1}M^{-1})\\
&=&P(\xi<h^{-1}(K(0)n^{-1}M^{-1}))\\
&=&h^{-1}(K(0)n^{-1}M^{-1}). \end{eqnarray*}
Hence, the series in equation (\ref{P}) is $\sum_{n=1}^\infty h^{-1}(K(0)n^{-1}M^{-1})$ and this converges or diverges, by monotonicity, with
$\int_0^\infty h^{-1}(K(0)t^{-1}M^{-1})dt$ and after change of variables we have $M^{-1}K(0)\int_0^\infty s^{-2}h^{-1}(s)ds$.\\
Now, integration by parts together with $h\geq aF$ shows that the latter integral converges and diverges with $\int_0^1(1/h)dF$.
\begin{eqnarray*}\int_0^\infty s^{-2}h^{-1}(s)ds&=&\int_0^1s^{-2}h^{-1}(s)ds+\int_1^\infty s^{-2}h^{-1}(s)ds\\
&\leq&\int_0^1s^{-2}h^{-1}(s)ds+\int_1^\infty s^{-2}\frac{s}{a}ds\\
&\leq&\left[s^{-2}\int h^{-1}(s)ds-\int (-2)s^{-3}\left(\int h^{-1}(s)ds\right)ds\right]_0^1+\frac{1}{a}\int_1^\infty\frac{1}{s}ds\\
&=&\left[s^{-2}\int h^{-1}(s)ds+2\int s^{-3}\left(\int h^{-1}(s)ds\right)ds\right]_0^1+\frac{1}{a}\int_1^\infty\frac{1}{s}ds
\end{eqnarray*}

Hence, $\int_0^1(1/h)dF=+\infty$ implies, by the divergence half of (\ref{P}), that $P(D_n\ i.o.)=1$ and therefore $P(B_n\ i.o.)=1$, for all $M>0$. Since, $M$ is arbitary.\\
Hence, (ii) is proved.
\begin{rem} If $\int_0^1(1/h)dF<\infty$ then $P(d_h(R_n,0)>nM\ i.o.)=0$ for all $M>0$. \end{rem}
We now prove (i)
Suppose, $h\in\mathcal{H}(\nearrow)$. Let $\epsilon>0$ and choose $\theta$ so small that $\int_0^\theta(1/h)dF<\epsilon/2$. Then
\begin{align}\label{A}d_h(F_{n,b}, F)\leq\sup_{0<t\leq\theta}\left(\frac{F_{n,b}(t)}{h(t)}\right)+\sup_{0<t\leq\theta}\left(\frac{t}{h(t)}\right)+\sup_{\theta\leq t\leq1}\frac{|F_{n,b}(t)-t|}{h(\theta)}\end{align}
\begin{eqnarray*}\sup_{0<t\leq\theta}\left(\frac{F_{n,b}(t)}{h(t)}\right)&=&\sup_{0<t\leq\theta}\left(\frac{\frac{1}{nb}\sum_{i=1}^n\int_{-\infty}^tk\left(\frac{x-\xi_i}{b}\right)dx}{h(t)}\right)\\
&\leq&\frac{\frac{1}{nb}\sum_{i=1}^n\int_{-\infty}^\theta k\left(\frac{x-\xi_i}{b}\right)dx}{h(\xi_i)}\\
&=&\frac{\frac{1}{n}\sum_{i=1}^nK\left(\frac{\theta-\xi_i}{b}\right)}{h(\xi_i)}\\
&\rightarrow&\int_0^\theta(1/h)dF\ w\ .\ p\ .\ 1\ by\ the\ ordinary\ strong\ law\ of\ large\ numbers. \end{eqnarray*}
$\because\frac{t}{h(t)}\leq\int_0^t(1/h)dF$ which implies
$\sup_{0<t\leq\theta}(t/h(t))\leq\int_0^\theta(1/h)dF$.
Now from equation (\ref{R}) we can say that the third term in equation (\ref{A}) converges to zero w. p. 1.\\
$\therefore$ We can write $$\limsup_{n\rightarrow\infty}d_h(F_{n,b}, F)<\epsilon/2+\epsilon/2=\epsilon\ w\ .\ p\ .\ 1\ for\ any\ \epsilon>0.$$
Hence, (i) is proved.

\textbf{Proof of Corollary 1.} From equation (\ref{L}) we can write that $d_h(F_{n,b},0)\rightarrow d_h(F,0)$ w.p.1 a $n\rightarrow\infty$. Hence, for any $\tau>1$, we can write $$P(d_h(F_{n,b},0)>\tau d_h(F,0)\ i.o.)=0.$$

\textbf{Proof of Theorem 2 considering distribution function $F_{n,b}$.} Note that it suffices to prove only the upper bound of (1) and (5): by replacing $\xi$ by $1-\xi_i$, by interchanging $\tau_1$ and $\tau_2$, and by use of symmetry about the identity function, the upper bound of (1) implies the remaining inequalities in (1) and (4): similarly (5) implies the remaining inequalities (2), (3) and (6). The proof of (5) is similar to the proof of (8) in Theorem 1 from Wellner \citep{Wellner2}.

 To prove the upper bound of (1), let $\alpha=1/\tau_1$ and $\tau>1$. We define $F_{n,b}^*=F_{n,b}-F$ and
$$E_n=\left\{\sup_{0<t\leq1}\frac{|F_{n,b}(t)|}{\tau t^\alpha}\geq1\right\}.$$
From Corollary 1.1 we can write that $P(E_n\ i.o.)=0$. Hence for $n\geq N(\omega, \alpha)$,
$$|F_{n,b}(t)|\leq\tau t^\alpha,\ 0\leq t\leq1$$
or,
$$F_{n,b}(t)\leq(1+\tau)t^\alpha,\ 0\leq t\leq1.$$
This implies that for $n\geq N(\omega, \alpha)$ and all $\omega$ in a set with probability one
$$F_{n,b}(t)\leq(t/\lambda)^\alpha,\ 0\leq t\leq1$$
where $0<\lambda\equiv2^{-1}(1+\tau)^{-\tau_1}<\frac{1}{2}$. Hence the upper bound of (1) is proved.

\textbf{Proof of Theorem 3.} Now, for $0\leq t\leq1$, we define $\psi_n(t)=-\int_t^1J_ndF$ so that $\frac{c_{ni}}{n}=\left[\psi_n\left(\frac{i}{n}\right)-\psi_n\left(\frac{(i-1)}{n}\right)\right]$ for $0\leq t\leq1$.
\begin{eqnarray*}T_n&=&\int_0^1g(F_{n,b}^{-1})J_ndF \\
&=&\sum_{i=1}^ng(\xi_{ni})\left[\psi_n\left(\frac{i}{n}\right)-\psi_n\left(\frac{(i-1)}{n}\right)\right] \\
&=&-\psi_n(0)g(\xi_{n1})-\sum_{i=1}^{n-1}\psi_n\left(\frac{i}{n}\right)[g(\xi_{ni+1})-g(\xi_{ni})] \\
&=&-\psi_n(0)g(\xi_{n1})-\int_{\xi_{n1}}^{\xi_{nn}}\psi_n(F_{n,b})dg\ a\ .\ s. \end{eqnarray*}
where the second integral representation uses the fact that it is a.s. true that no $\xi_{ni}$ takes on one of the countable number of values at which $g$ is discontinuous.\\
Now,
\begin{eqnarray*}\mu_n&=&\int_0^1J_ngdF \\
&=&\int_{\xi_{n1}}^{\xi_{nn}}gd\psi_n+\int_{[\xi_{n1},\xi_{nn}]^c}J_ngdF \\
&=&-\int_{\xi_{n1}}^{\xi_{nn}}\psi_ndg+\psi_n(\xi_{nn})g(\xi_{nn})-\psi_n(\xi_{n1})g(\xi_{n1})+\int_{[\xi_{n1},\xi_{nn}]^c}J_ngdF\ a\ .\ s. \end{eqnarray*}
$\therefore\ T_n-\mu_n=-(A_{n1}+A_{n2}+A_{n3}+A_{n4})$, where $$A_{n1}=\int_{\xi_{n1}}^{\xi_{nn}}Z_{n}(F_{n,b}-F)dg=\int_0^1Z_n^*(F_{n,b}-F)dg$$ with $Z_n=\frac{(\psi(F_{n,b})-\psi)}{(F_{n,b}-F)}$ where $Z_n^*$ is equal to $Z_n$ on [$\xi_{n1},\xi_{nn}$) and is equal to 0 otherwise, and
\begin{eqnarray*}A_{n2}&=&g(\xi_{n1})[\psi_n(0)-\psi_n(\xi_{n1})] \\
                 A_{n3}&=&g(\xi_{nn})\psi_n(\xi_{nn}) \\
                 A_{n4}&=&\int_{[\xi_{n1},\xi_{n2}]^c}gJ_ndF \end{eqnarray*}
Now we need to show $A_{n2}$, $A_{n3}$ and $A_{n4}$ are negligible. From the proof of Theorem 1 of Shorack \cite{shorack72} we can say that $A_{n2}$, $A_{n3}$ and $A_{n4}$ are negligible. So we can write $(A_{n2}+A_{n3}+A_{A4})\rightarrow0$ w.p.1 as $n\rightarrow\infty$.
Then our aim is to show that $A_{n1}\rightarrow0$ w. p. 1 as $n\rightarrow\infty$.\\
Now, by Assumption 3, when $b_1,\ b_2>0$ we have
$$|Z_n|=\left|\frac{\int_F^{F_{n,b}}J_ndF}{F_{n,b}-F}\right|\leq\frac{\int_F^{F_{n,b}}BdF}{F_{n,b}-F}\leq B\vee B(F_{n,b}).$$
Now we choose $\tau_1$, $\tau_2$ in Theorem 2 so that $b_1\tau_1=b_1+\delta/4$, $b_2\tau_2=b_2+\delta/4$, and fix $\omega\in A$. Then, for $n\geq N_\omega$, (2) and (3) imply that
\begin{eqnarray}\label{Z}|Z_n^*|&\leq&M_{1,2}MF^{-b_1\tau_1}(1-F)^{-b_2\tau_2} \notag  \\
                &=&M_{1,2}MF^{-(b_1+\delta/4)}(1-F)^{-(b_2+\delta/4)}\notag \\
                &=&M_{1,2}B[F(1-F)]^{\delta/4} \end{eqnarray}
for some constant $M_{1,2}$ depending on $\beta$ of Theorem 2. Clearly we can say that equation (0.7) holds if either $b_1$ or $b_2$ equals zero. If $b_1\ or\ b_2<0$ then by use of (1) of Theorem 2 and an argument similar to that given for $b_1\ or b_2>0$ also yields equation (0.7).
Now, w.p.1, for $n\geq N_\omega$
\begin{eqnarray*}|A_{n1}|&\leq&\int_0^1|Z_n^*||F_{n,b}-F|d|g|, \\
                &\leq&M_{1,2}\int_0^1B[F(1-F)]^{\delta/4}(|F_{n,b}-F|/h^*)h^*d|g|,\ using\ (\ref{Z}) \\
                &\leq&M_{1,2}\int_0^1Bh(|F_{n,b}-F|/h^*)d|g|,\ since\ h^*[F(1-F)]^{\delta/4}=h \\
                &\leq&M_{1,2}d_{h^*}(F_{n,b}-F,0)\int_0^1Bhd|g|.
\end{eqnarray*}
As $h^*\in\mathcal{H}$, so Theorem 1 implies that $d_{h^*}(F_{n,b}-F,0)\rightarrow 0$ w.p.1 as $n\rightarrow\infty$. Also, $\int_0^1Bhd|g|<\infty$ by Assumption 3. Hence, $A_{n1}\rightarrow0$ w.p.1 as $n\rightarrow\infty$.\\
$\therefore$ We can write $\lim_{n\rightarrow\infty}(T_n-\mu_n)=0\ w.p.1.$\\

\textbf{Proof of Corollary 2.} The proof is similar to the proof of Corollary 2 of Wellner \citep{Wellner1}.\

\textbf{Proof of Theorem 4.} If we show that $\lim_{n\rightarrow\infty}\mu_n=\mu$, then Corollary 3.1 with $\mu_{\infty}=\mu$ is in force and the proof is complete. But, by Assumption 3 we have $|J_ng|\leq M^2[F(1-F)]^{-1+\delta}$ which is in $L^1(F)$. Again from Assumption 4 we have $J_n(t)g(t)\rightarrow J(t)g(t)$ for all $t\in (0,\ 1)$. Therefore, by the dominated convergence theorem, we can write
$$\mu_n=\int_0^1J_ngdF\rightarrow \int_0^1JgdF=\mu.$$
Hence, $$\lim_{n\rightarrow\infty}T_n=\mu\ w.p.1.$$

\textbf{Proof of Theorem 5.} Now, for $0\leq t\leq1$, we define $\psi_n(t)=-\int_t^1J_ndF$ so that $\frac{c_{ni}}{n}=\left[\psi_n\left(\frac{i}{n}\right)-\psi_n\left(\frac{(i-1)}{n}\right)\right]$ for $0\leq t\leq1$. \begin{eqnarray*}T_n-\mu_n&=&\int_0^1g(F_{n,b}^{-1}(t))d\psi_n(t)-\int_0^1g(t)d\psi_n(t) \\
&=& \int_0^1g(t)d[\psi_n(F_{n,b}(t))-\psi_n(t)]
\end{eqnarray*}
Integrating by parts, we have
\begin{align} \label{S1}T_n-\mu_n=\lim_{\theta\rightarrow0}[g(t)\{\psi_n(F_{n,b}(t))-\psi_n(t)\}]_{\theta}^{1-\theta}
-\int_0^1[\psi_n(F_{n,b}(t))-\psi_n(t)]dg(t) \end{align}
Now, for $0<t<\xi_{n1}$, we have
\begin{align*} |g(t)\{\psi_n(F_{n,b}(t))-\psi_n(t)\}|\leq D(t)\int_0^tB(u)du\leq Mt^\delta\rightarrow0\ as\ t\rightarrow0\end{align*}
 A similar argument for $\xi_{nn}<t<1$ holds. Thus equation (\ref{S1}) can be written as
\begin{align*} T_n-\mu_n=-\int_0^1[\psi_n(F_{n,b}(t))-\psi_n(t)]dg(t) \end{align*}
Now, we can write the above equation as
\begin{align}\label{S2}T_n-\mu_n=-\gamma_n-S_n,\end{align}
where $\gamma_n=\int_0^1[\psi_n(F_{n,b}(t))-\psi_n(t)-\{F_{n,b}(t)-t\}J(t)]dg(t)$ and
$S_n=\int_0^1[F_{n,b}(t)-t]J(t)dg(t)$.\

 Now, \begin{align}\label{N}\sqrt{n}\gamma_n=\int_0^1U_n(t)A_n(t)dg(t),\end{align}
where $U_n(t)=\sqrt{n}(F_{n,b}(t)-t)$ and $A_n(t)=\frac{1}{F_{n,b}(t)-t}\int_t^{F_{n,b}(t)}J_n(u)du-J(t)$.\

 Now, $$\sqrt{n}|\gamma_n|\leq ||U_n/h||\int_0^1|A_n(t)|h(t)d|g|(t),$$
where $||\cdot||$ denotes the sup-norm on ($0, 1$).
Now, for $\xi_{n1}\leq t\leq\xi_{nn}$, it follows from Assumption 3 that
\begin{eqnarray*}|A_n(t)|&\leq &\frac{1}{F_{n,b}(t)-t}\int_t^{F_{n,b}(t)}|J_n(u)|du+|J(t)| \\
&\leq & B(t)\vee B(F_{n,b}(t))+B(t)
\end{eqnarray*}
From Theorem 2, $\exists$ a set $A\subset\Omega$ with $P(A)=1$ for a significantly large $n$. So for $\xi_{n1}\leq t<\xi_{nn}$ we have
$$|A_n(t)|h(t)\leq Mt^{1-\delta/2-b_1(1+\tau_1)}(1-t)^{1/2-\delta/2-b_2(1+\tau_2)}.$$

Now for $0<t<\xi_{n1}$ and $F_{n,b}(t)=0$, we have
$$|A_n(t)|\leq\frac{1}{t}\int_0^1B(u)du+B(t)\leq Mt^{-b_1}.$$

Similarly, we have for $\xi_{nn}\leq t<1.$

$\therefore$ On a set $A$ we have $$|A_n(t)|h(t)\leq Mt^{1-\delta/2-b_1(1+\tau_1)}(1-t)^{1/2-\delta/2-b_2(1+\tau_2)}.$$
Now, from Assumption 3 we see that the right-hand side of equation (\ref{N}) is $|g|$-integrable and by Assumption 4 we have $A_n(t)\rightarrow0$, $|g|$-a.e as $n\rightarrow\infty$ with probability one.\

Hence we can write $\sqrt{n}|\gamma_n|\rightarrow0$ as $n\rightarrow\infty$ with probability one.\

$\therefore$ Equation (\ref{S2}) can be written as
\begin{align} \label{F}\sqrt{n}(T_n-\mu_n)=-\sqrt{n}S_n. \end{align}
Now, $$\sqrt{n}S_n=\int_0^1\mathbb{V}_n(t)J(t)dg(t),$$ where $\mathbb{V}_n(t)=\sqrt{n}(F_{n,b}(t)-t)$.
We now define $$S=\int_0^1\mathbb{V}(t)J(t)dg(t)$$ so that $S$ is a $N(0, \sigma^2)$ random variable and $\sigma^2$ is finite by Assumption 3.\

Now using Corollary 3 and dominated convergence theorem, we can write that
\begin{align*}\sqrt{n}S_n=\int_0^1\mathbb{V}_n(t)J(T)dg(t)\rightarrow S=\int_0^1\mathbb{V}(t)J(t)dg(t) \end{align*}
And hence we can write $$\sqrt{n}(T_n-\mu_n)\rightarrow -\int_0^1\mathbb{V}(t)J(t)dg(t).$$
\newpage
\bibliographystyle{Chicago}
\bibliography{refs}
\begin{verbatim}https://www.macrotrends.net\end{verbatim}
\begin{verbatim}https://www.nseindia.com\end{verbatim}
\newpage
\begin{table}[ht]
\caption{Ratios estimated using different estimators of distortion risk measures.}
 \centering
 \begin{tabular}{|l|l|l|l|l|l|l|l|}
\hline
Ratio & $\beta$ & n & GPD & Student t & N(0,1) & GARCH \\ [0.5ex]
&&&&&&$\alpha_1=0.061$\\
&&&&&&$\beta_1=0.932$\\
 \hline
 & $10$ &$30$&\space $0.953$ & $0.958$ & $0.972$ & $0.9745$\\
     &&$100$&\space  $0.977$ & $0.982$ & $0.989$ &$0.990$\\
      &&$250$&\space $0.993$ & $0.994$ & $0.996$ &$0.996$\\

$\frac{MSE2}{MSE1}$& $5$ &$30$&\space $0.790$ & $0.810$ & $0.905$& $0.943$  \\
     &&$100$&\space  $0.937$ & $0.956$ & $0.976$ & $0.978$ \\
      &&$250$&\space $0.986$ & $0.985$ & $0.991$ &$0.991$\\

    &$1$& $30$ &\space $0.704$ & $0.638$ & $0.758$ &$0.679$\\
    &&  $100$&\space $0.885$ & $0.878$ & $0.848$ & $0.880$\\
    && $250$ &\space $0.968$ & $0.955$ & $0.939$ & $0.949$ \\ [1ex]
      \hline
 \end{tabular}\label{tab1}
\end{table}
\clearpage
\newpage
\addtolength{\oddsidemargin}{-.7in}%
\addtolength{\evensidemargin}{-.7in}%
\addtolength{\textwidth}{1.4in}%
\begin{table}[ht]
\centering
\begin{threeparttable}
\caption{Estimates of exponential spectral risk measure of future index  and the standard deviation.}
\begin{small}
 \begin{tabular}{|l|l|l|l|l|l|l|}
\hline
$\shortstack{Future \\Index}$ & $\beta=1$ & $\beta=5$ & $\beta=10$ & $\beta=20$ & $\beta=100$ & $\beta=200$ \\
 \hline
 \multicolumn{7}{|c|}{Time period $1/01/1991-31/12/2003$}\\
 \hline
$\shortstack{Nikkei\\$225$}$ & $-0.218(0.0538)$ & $-0.946(0.0849)$ & $-1.499(0.1284)$ & $-2.28(0.2018)$ & $-6.79(0.488)$ & $-11.84(0.5763)$  \\
DAX & $-0.230(0.0471)$ & $-0.918(0.0792)$ & $-1.456(0.1169)$ & $-2.23(0.1764)$ & $-6.79(0.383)$ & $-11.88(0.455)$  \\
$\shortstack{FTSE\\$100$}$ & $-0.183(0.0384)$ & $-0.748(0.0664)$ & $-1.212(0.1069)$ & $-1.91(0.1596)$ & $-6.26(0.384)$ & $-11.26(0.508)$  \\
$\shortstack{Hang \\Seng}$ & $-0.253(0.058)$ & $-0.998(0.099)$ & $-1.573(0.151)$ & $-2.39(0.24)$ & $-7.01(0.63)$ & $-12.11(0.78)$  \\ [1ex]
      \hline
       \multicolumn{7}{|c|}{Time period $2/01/2004-31/12/2008$}\\
       \hline
       $\shortstack{Future \\Index}$ & $\beta=1$ & $\beta=5$ & $\beta=10$ & $\beta=20$ & $\beta=100$ & $\beta=200$ \\
 \hline
$\shortstack{Nikkei\\$225$}$ & $-0.241(0.0634)$ & $-1.0274(0.0900)$ & $-1.6727(0.1454)$ & $-2.604(0.1900)$ & $-7.705(0.3990)$ & $-12.984(0.4870)$  \\
DAX & $-0.186(0.0400)$ & $-0.837(0.0650)$ & $-1.392(0.1023)$ & $-2.222(0.1654)$ & $-7.062(0.2983)$ & $-12.248(0.4201)$  \\
$\shortstack{FTSE\\$100$}$ & $-0.179(0.0298)$ & $-0.788(0.0598)$ & $-1.321(0.0989)$ & $-2.138(0.1456)$ & $-6.911(0.2778)$ & $-12.087(0.4986)$  \\
$\shortstack{Hang \\Seng}$ & $-0.223(0.0547)$ & $-0.992(0.0965)$ & $-1.636(0.1460)$ & $-2.553(0.1990)$ & $-7.402(0.5494)$ & $-12.551(0.6549)$  \\
\hline
      \multicolumn{7}{|c|}{Time period $2/01/2009-2/01/2019$}\\
       \hline
Nikkei$225$ & $-0.217(0.059)$ & $-0.856(0.087)$ & $-1.357(0.1375)$ & $-2.09(0.2301)$ & $-6.51(0.501)$ & $-11.55(0.6040)$  \\
DAX & $-0.216(0.0510)$ & $-0.832(0.0801)$ & $-1.328(0.1232)$ & $-2.06(0.1890)$ & $-6.48(0.3944)$ & $-11.53(0.4956)$  \\
FTSE$100$ & $-0.168(0.0599)$ & $-0.695(0.0794)$ & $-1.135(0.1134)$ & $-1.81(0.1794)$ & $-6.13(0.4576)$ & $-11.12(0.5909)$  \\
Hang Seng & $-0.184(0.0621)$ & $-0.783(0.1267)$ & $-1.216(0.1898)$ & $-1.94(0.3235)$ & $-6.22(0.7102)$ & $-11.19(0.8321)$  \\ [1ex]
      \hline
 \end{tabular}
 \end{small}
 \begin{tablenotes}
      \small
      \item Notes: Estimates are in daily $\%$ return.
    \end{tablenotes}
  \end{threeparttable}
\label{table1:VaR}
\end{table}

\begin{table}[ht]
\caption{$90\%$ confidence intervals.}
 \centering
 \begin{small}
 \begin{tabular}{|l|l|l|l|l|l|l|}
\hline
\multicolumn{7}{|c|}{Time period $1/01/1991-31/12/2003$}\\
\hline
$\shortstack{Future \\Index}$ & $\beta=1$ & $\beta=5$ & $\beta=10$ & $\beta=20$ & $\beta=100$ & $\beta=200$  \\ [0.5ex]
 \hline
$\shortstack{Nikkei\\$225$}$ & [-0.0024 -0.0020] & [-0.0097 -0.0093] & [-0.0152 -0.0148] & [-0.0230 -0.0226] & [-0.0681 -0.0677] & [-0.1186 -0.1182] \\
DAX & [-0.0025 -0.0021] & [-0.0094 -0.0090] & [-0.0148 -0.0144] & [-0.0225 -0.0221] & [-0.0681 -0.0677] & [-0.1190 -0.1186]  \\
$\shortstack{Hang \\Seng}$ & [-0.0027 -0.0023] & [-0.0102 -0.0098] & [-0.0159 -0.0155] & [-0.0241 0.0237] & [-0.0703 -0.0699] & [-0.1213 -0.1209]  \\
$\shortstack{FTSE\\$100$}$ & [-0.0019 -0.0017] & [-0.0076 -0.0073] & [-0.0123 -0.0119] & [-0.0192 -0.0190] & [-0.0627 -0.0625] & [-0.1127 -0.1125]  \\ [1ex]
           \hline
       \multicolumn{7}{|c|}{Time period $2/01/2004-31/12/2008$}\\
       \hline
  $\shortstack{Future \\Index}$ &$\beta=1$ &$\beta=5$ &$\beta=10$ & $\beta=20$ & $\beta=100$ & $\beta=200$  \\ [0.5ex]
 \hline
$\shortstack{ Nikkei\\225}$ & [-0.0027 -0.0022] &[-0.0105 -0.0100] & [-0.0170 -0.0165] & [-0.0263 -0.0258] & [-0.0773 -0.0768] & [-0.1301 -0.1296] \\
DAX & [-0.0021 -0.0017] & [-0.0086 -0.0082] & [-0.0141 -0.0137] & [-0.0224 -0.0220] & [-0.0708 -0.0704] & [-0.1227 -0.1223]  \\
$\shortstack{Hang \\Seng}$ & [-0.0025 -0.0020] & [-0.0102 -0.0097] & [-0.0166 -0.0161] & [-0.0258 -0.0253] & [-0.0743 -0.0738] & [-0.1258 -0.1253]  \\
$\shortstack{FTSE\\$100$}$ & [-0.0020 -0.0016] & [-0.0081 -0.0077] & [-0.0134 -0.0130] & [-0.0216 -0.0212] & [-0.0693 -0.0689] & [-0.1210 -0.1207]  \\ [1ex]
      \hline
            \multicolumn{7}{|c|}{Time period $2/01/2009-2/01/2019$}\\
       \hline
       $\shortstack{Future \\Index}$ &$\beta=1$ &$\beta=5$ &$\beta=10$ & $\beta=20$ & $\beta=100$ & $\beta=200$  \\ [0.5ex]
 \hline
$\shortstack{ Nikkei\\225}$ & [-0.0024 -0.0020] &[-0.0088 -0.0084] & [-0.0138 -0.0134] & [-0.0211 -0.0207] & [-0.0653 -0.0649] & [-0.1157 -0.1153] \\
DAX & [-0.0023 -0.0020] & [-0.0085 -0.0081] & [-0.0135 -0.0131] & [-0.0207 -0.0204] & [-0.0650 -0.0646] & [-0.1155 -0.1151]  \\
$\shortstack{Hang \\Seng}$ & [-0.0020 -0.0017] & [-0.0080 -0.0077] & [-0.0123 -0.0120] & [-0.0196 -0.0192] & [-0.0624 -0.0620] & [-0.1121 -0.1117]  \\
$\shortstack{FTSE\\$100$}$ & [-0.0018 -0.0015] & [-0.0071 -0.0068] & [-0.0115 -0.0112] & [-0.0182 -0.0179] & [-0.0614 -0.0612] & [-0.1157 -0.1153]  \\ [1ex]
      \hline
 \end{tabular}
 \end{small}
\label{table1:VaR}
\end{table}

\end{document}